A model rocket is launched from the ground with an inital velocity of 68 m/s. The function h = -4.9t + 68t models the height of the rocket in metres after t seconds. At what time is the rocket above 72 m above the ground?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

-4.9t + 68t = h
63.1t = 72
t = 72/63.1 = 1.14 seconds

Answer by ikleyn(52781)

.
A model rocket is launched from the ground with an inital velocity of 68 m/s. The function h = -4.9t + 68t models the height
of the rocket in metres after t seconds. At what time is the rocket above 72 m above the ground?
~~~~~~~~~~~~~~~~~~~~~~~

You  INCORRECTLY  presented your function.

The correct form is h = -4.9t^2 + 68t.     It is the standard form when the unit of length/distance  the"foot"  is used.

To answer the question, you need to solve the quadratic equation
-4.9*t^2 + 68t = 72.

-------------------
To see many other similar solved problems, look into the lessons
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


===============
Be aware ! The solution by  @addingup  is   I N C O R R E C T  !.